# SGD with Clipping is Secretly Estimating the Median Gradient

### v1

@@ B.3 Approximations based on the Stochastic Proximal Point: General $\mathcal{D}$
$\mathrm{prox}_{\tau\phi}(\bm{m}_{0})=\underset{\bm{m}\in\mathbb{R}^{d}}{\rm{argmin}}~{}\phi(\bm{m})+\frac{1}{2\tau}\|\bm{m}-\bm{m}_{0}\|^{2}.$
@@
$\begin{cases}\text{Sample $\bm{g}_{t}$ i.i.d. from $\mathcal{G}$}\\ \bm{m}_{t+1}=\mathrm{prox}_{\tau\mathcal{D}(\cdot-\bm{g}_{t})}(\bm{m}_{t}).\end{cases}$ (28)
@@
Let $\mathcal{D}$ verify Assumption B.1. The SPP update (28) for solving (8) can be equivalently written as
@@
where $\mathcal{D}^{*}(\bm{m})=\sup_{\bm{y}}\left<\bm{y},\bm{m}\right>-\mathcal{D}(\bm{y})$ is the Fenchel conjugate of $\mathcal{D}$, and $\bm{g}_{t}$ is sampled i.i.d. from $\mathcal{G}$ at each iteration.
@@
Here we make use of proximal calculus rules, which can found in most textbooks on convex analysis, such as chapter 6 in (Beck, 2017). Starting from (9), with $\bm{y}=\bar{\bm{y}}+\bm{g}_{t}$, we have
@@
$\mathrm{prox}_{\tau\mathcal{D}}(\bm{m})=\bm{m}-\mathrm{prox}_{\tau\mathcal{D}^{*}(\frac{\cdot}{\tau})}(\bm{m})$
in (29) gives
@@
B.4 Approximations based on the Stochastic Proximal Point: Particular Cases of $\mathcal{D}$
$\ell_{p}$-norm.
We now specialize the previous results for when $\mathcal{D}$ is an $\ell_{p}$-norm. See 4.1
@@ B.4 Approximations based on the Stochastic Proximal Point: Particular Cases of $\mathcal{D}$
see Example 6.47 in (Beck, 2017). Using this in (29) gives
@@
$\ell_{1}$-norm.
See 4.3
@@
See 4.2
@@
Follows directly from (10), by using that
Huber function.
The Huber function is given by
$H_{\mu}:\mathbb{R}^{d}\to\mathbb{R},\quad H_{\mu}(\bm{z})=\begin{cases}\frac{1}{2\mu}\|\bm{z}\|^{2}&\quad\|\bm{z}\|\leq\mu,\\ \quad\|\bm{z}\|-\frac{\mu}{2}&\quad\text{else}.\end{cases}$
It operates like the squared loss for arguments with small norm, and like the $\ell_{2}$-norm for arguments with large norms (potentially outliers).
Lemma B.4.
Update (9) with $\mathcal{D}=H_{\mu}$ is given by
$\bm{m}_{t+1}$ $=\bm{g}_{t}+\big{(}1-\tfrac{\tau}{\max\{\|\bm{m}_{t}-\bm{g}_{t}\|,\tau+\mu\}}\big{)}(\bm{m}_{t}-\bm{g}_{t})=\big{(}1-\tfrac{\tau}{\max\{\|\bm{m}_{t}-\bm{g}_{t}\|,\tau+\mu\}}\big{)}\bm{m}_{t}+\tfrac{\tau}{\max\{\|\bm{m}_{t}-\bm{g}_{t}\|,\tau+\mu\}}\bm{g}_{t}.$ (32)
Proof.
Follows from (9) and the fact that the proximal operator of the Huber function (see (Beck, 2017, Example 6.66)) is given by $\mathrm{prox}_{\tau H_{\mu}}(\bm{z})=\Big{(}1-\frac{\tau}{\max\{\|\bm{z}\|,\tau+\mu\}}\Big{)}\bm{z}.$
The estimator (32) has the advantage that it is a relaxation of vectorwise clipping (12), as $01-\frac{\tau}{\max\{\|\bm{m}_{t}-\bm{g}_{t}\|,\tau+\mu\}}\geq 1-\frac{\tau}{\tau+\mu}>$ 0. However, it comes with the disadvantage that it has two hyperparameters $\tau,~{}\mu$ instead of one.
B.5 Proof of Proposition 4.4: Convergence when $\mathcal{D}$ is a Norm
See 4.4
Proof.
Since we assume that $\mathcal{D}$ is a norm over $\mathbb{R}^{d}$ then it convex and continuous. Moreover, since all norms are equivalent in $\mathbb{R}^{d}$, we have that $\mathcal{D}(\bm{w})\leq C\|\bm{w}\|_{2}$ for some constant $G>0$. Thus $\mathcal{D}$ is a $G$-Lipschitz function. If now we write $\varphi_{\bm{g}}:=\mathcal{D}(\cdot-\bm{g})$, it is clear that it is convex and Lipschitz. Therefore, $\varphi=\mathbb{E}_{\bm{g}}\left[\varphi_{\bm{g}}\right]$ is itself a convex $G$-Lipschitz function. We can now conclude with Theorem 4.4 from (Davis & Drusvyatskiy, 2019). ∎
@@
In the context of Proposition 4.4, assume that $\mathcal{D}=\|\cdot\|_{1}$ and that $\mathcal{G}$ is a sum of Dirac measures $\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\bm{g}^{(i)}}$. Then we have further
@@
To obtain the desired bound, and given that we already have the conclusion of Proposition 4.4, it is enough for us to show that there exists $\mu>0$ such that
$(\forall\bm{m}\in\mathbb{R}^{d})\quad\mu\operatorname{dist}(\bm{m},{\rm{argmin}}~{}\varphi)\leq\varphi(\bm{m})-\inf\varphi.$
It is a standard result from variational analysis that this property [^5] It is sometimes referred to as (weak) sharp minima (Ferris, 1991; Burke & Ferris, 1993), (superlinear)(linear) conditioning (Lemaire, 1992; Cornejo et al., 1997; Lemaire, 1998), or error bound (Lewis & Pang, 1998). is true if and only if the following inequality is satisfied (see Theorem 5.1 in (Cornejo et al., 1997) or Proposition 3.1 in (Lemaire, 1998))
$(\forall\bm{m}\in\operatorname{dom}\partial\varphi)\quad\|\partial\varphi(\bm{m})\|_{-}\geq\mu.$
Our goal now is to prove the above inequality, from which the conclusion will follow. Let us consider $\bm{m}\notin{\rm{argmin}}~{}\varphi$, and use standard calculus to write
$\partial\varphi(\bm{m})=\frac{1}{n}\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}-{\bm{g}}_{i})=\frac{1}{n}\sum_{i=1}^{n}\big{(}\operatorname{sgn}(\bm{m}_{j}-({\bm{g}}_{i})_{j})\big{)}_{j=1,\ldots,d}=\frac{1}{n}\Big{(}\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}_{j}-({\bm{g}}_{i})_{j})\Big{)}_{j=1,\ldots,d}.$
Our goal is to show that $\|\partial\varphi(\bm{m})\|_{-}\geq\frac{1}{n}$, which is equivalent to show that $\|\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}-{\bm{g}}_{i})\|_{-}\geq 1$. Because $\bm{m}\notin{\rm{argmin}}~{}\varphi$, we know that $0\notin\partial\varphi(\bm{m})$ so there must exist a coordinate $j$ such that the $j$-th coordinate of $\partial\varphi(\bm{m})$ does not contain zero, in other words such that $0\notin\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}_{j}-({\bm{g}}_{i})_{j})$. Using the fact that $\|\cdot\|_{2}\geq\|\cdot\|_{\infty}$, we can therefore lower bound $\|\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}-{\bm{g}}_{i})\|_{-}\geq|\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}_{j}-({\bm{g}}_{i})_{j})|_{-}$. Let us denote $s_{i}:=\operatorname{sgn}(\bm{m}_{j}-({\bm{g}}_{i})_{j})$, so that we want to prove that $|\sum_{i=1}^{n}s_{i}|_{-}\geq 1$. Let us now consider a few cases.
@@
    where the latter is the distance from the number $s$ to the interval $[-k,+k]$. Now remember that $0\notin\sum_{i=1}^{n}s_{i}$, so this distance cannot be zero, which means that $|s|>k$. Moreover both $s$ and $k$ are rational numbers, so this distance must be greater or equal to $1$. This concludes the proof.
@@ Appendix C Supplementary Information on Experiments
For easier interpretation, we use the first order approximation of momentum (14), namely $\bm{m}_{t+1}=\left(1-\tau\right)\bm{m}_{t}+\tau\bm{g}_{t}$. That is, we replace $\frac{\tau}{1+\tau}$ with its first-order Taylor approximation around zero, which is equal to $\tau$. Note that $\tau$ is typically small, for example for $\tau=0.01$ we have $\frac{\tau}{1+\tau}\approx 0.0099$. This applies only to the experiments with estimating a fixed vector (or gradient) in Section 6.1.
For the language modeling experiments, all details are identical to (Kunstner et al., 2023), Section A.1. They also provide implementations for all tasks at https://github.com/fKunstner/noise-sgd-adam-sign, which we use. For the training runs for language modeling, we use batch size $256$ for PTB, $320$ for WikiText-2, and $32$ for SQuAD.
@@ C.1 Additional Plots
Here, we show the results for estimating the full gradient of a transformer architecture, with weights fixed at initialization. Fig. 7 shows the results for PTB where we tried three different batch sizes $\{64,256,1024\}$. Fig. 8 shows the results for WikiText-2.
Discussion. From Fig. 7, we observe two phenomena: in the long run, momentum attains the lowest error for estimating the full batch gradient. However, the initial decrease of the error is much faster for CClip, followed by VClip. This is important when using these estimates within a training setup such as (17), where we only do one iteration of gradient estimation, followed by an update of the weight (and hence a change in the full-batch gradient). Secondly, we observe that the difference in convergence speeds is most pronounced when the batch size increases. Hence, being robust to outliers seems not to be solvable only by increasing the batch size and thus decreasing the noise of the mini-batch gradient. In fact, the contrary seems to be the case. This is similar to the observations made in (Kunstner et al., 2023).

### v2

@@ B.2 Approximations Based on the Stochastic Proximal Point: General $\mathcal{D}$
$\mathrm{prox}_{\tau\phi}(\bm{m}_{0})=\underset{\bm{m}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\phi(\bm{m})+\frac{1}{2\tau}\|\bm{m}-\bm{m}_{0}\|^{2}.$
@@
$\begin{cases}\text{Sample $\bm{g}_{t}$ i.i.d. from $\mathcal{G}$}\\ \bm{m}_{t+1}=\underset{\bm{m}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\mathcal{D}(\bm{m}-\bm{g}_{t})+\frac{1}{2\tau}\|\bm{m}-\bm{m}_{t}\|^{2}.\end{cases}$ (27)
@@
Let $\mathcal{D}$ verify Assumption˜B.1. The SPP update (27) for solving (9) can be equivalently written as
@@
where $\mathcal{D}^{*}(\bm{m}):=\sup_{\bm{y}}\big(\left<\bm{y},\bm{m}\right>-\mathcal{D}(\bm{y})\big)$ is the Fenchel conjugate of $\mathcal{D}$, and $\bm{g}_{t}$ is sampled i.i.d. from $\mathcal{G}$ at each iteration.
@@
Here we make use of proximal calculus rules, which can found in most textbooks on convex analysis, for example Chapter 6 in Beck (2017). Starting from (27), and applying the variable transformation $\bm{y}=\bar{\bm{y}}+\bm{g}_{t}$, we have
@@
$\mathrm{prox}_{\tau\mathcal{D}}(\bm{m})=\bm{m}-\mathrm{prox}_{\tau\mathcal{D}^{*}(\frac{\cdot}{\tau})}(\bm{m}).$
Plugging into (28) we have
@@
B.3 Approximations Based on the Stochastic Proximal Point: Particular Cases of $\mathcal{D}$
We first characterize the SPP update (9) for when $\mathcal{D}$ is the $\ell_{p}$-norm, and show it simply require to project the sampled gradient onto a certain ball of radius $\tau$ centered at $\bm{m}_{t}$.
Proposition B.4.
$=\mathrm{Proj}_{\mathbb{B}_{q}(\bm{m}_{t},\tau)}\big(\bm{g}_{t}\big).$ (32)
@@ B.3 Approximations Based on the Stochastic Proximal Point: Particular Cases of $\mathcal{D}$
see Example 6.47 in Beck (2017). Using this in (28) gives
@@
Using Proposition˜B.4 we can now develop closed form updates of the SPP method for when $\mathcal{D}$ is the $\ell_{1}$- or the $\ell_{2}$-norm, with both methods being related to clipping.
See 3.2
@@
We end this section with the closed form updates of the SPP method for when $\mathcal{D}$ is the Huber function.
See 3.3
Proof.
Follows from (9) and the fact that the proximal operator of the Huber function is given by $\mathrm{prox}_{\tau H_{\mu}}(\bm{z})=\big(1-\frac{\mu\tau}{\max\{\|\bm{z}\|,\mu(1+\tau)\}}\big)\bm{z}.$ For a proof, see Beck (2017, Example 6.66), and note that here we have the additional factor $\mu$ in the definition of $H_{\mu}$. ∎
B.4 Approximations Based on Subgradient Updates
Here, $\tau>0$ is the learning rate, and $\bm{u}_{t}$ is a subgradient (of the convex subdifferential). Since $\mathcal{D}$ could be a non-differentiable function such as the $\ell_{2}$- or $\ell_{1}$-norm, we use subgradients instead of gradients.
Squared $\ell_{2}$-norm.
This is again (heavy-ball) momentum (cf. SGD-M), now with coefficient $\beta=1-\tau$. Note that in contrast to SPP, the coefficient $\beta$ could be negative, and is only in $[0,1)$ (as typically is the case for momentum) if $\tau\in[0,1]$.
$\ell_{1}$-norm.
If we choose $\mathcal{D}=\|\cdot\|_{1}$, update (33) gives
$\bm{m}_{t+1}=\bm{m}_{t}+\tau\operatorname{sgn}(\bm{g}_{t}-\bm{m}_{t}),$
where the $\operatorname{sgn}$-operator can take any value in [1,1][-1,1] for coordinates $i\in[d]$ such that $(\bm{g}_{t}-\bm{m}_{t})_{i}=0$. This recovers a version of sign-SGD (Riedmiller & Braun, 1993; Bernstein et al., 2018), where the $\mathrm{sign}$-operation is applied to the increment $\bm{m}_{t}-\bm{g}_{t}$. For example, if we reset $\bm{m}_{t}=0$ in every iteration, we obtain exactly sign-SGD.
@@ B.4 Approximations Based on Subgradient Updates
B.5 Rates for SPP
Now that we have presented how to compute the iterates (9) in practice, let us state its convergence properties. Since all these methods are instantiates of the SPP method, we can use standard results such as Davis & Drusvyatskiy (2019, Thm. 4.4).
Proposition B.5.
Let $\mathcal{D}$ be a norm, let $\phi(\bm{m}):=\mathbb{E}_{\bm{g}\sim\mathcal{G}}\left[\mathcal{D}(\bm{m}-\bm{g})\right]$, and let $\bm{m}_{*}\in\underset{\bm{m}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\phi(\bm{m})$. If $\bm{m}_{t}$ is generated by (9) then for $\bar{\bm{m}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\bm{m}_{t}$ we have
$\mathbb{E}\left[\phi(\bar{\bm{m}}_{T})-\phi(\bm{m}_{*})\right]=O\left(\frac{\|\bm{m}_{0}-\bm{m}_{*}\|^{2}}{\tau T}+\tau\right).$
@@ B.5 Rates for SPP
Since we assume that $\mathcal{D}$ is a norm over $\mathbb{R}^{d}$ then it convex and continuous. Moreover, since all norms are equivalent in $\mathbb{R}^{d}$, we have that $\mathcal{D}(\bm{w})\leq C\|\bm{w}\|_{2}$ for some constant $G>0$. Thus $\mathcal{D}$ is a $G$-Lipschitz function. If now we write $\phi_{\bm{g}}:=\mathcal{D}(\cdot-\bm{g})$, it is clear that it is convex and Lipschitz. Therefore, $\phi=\mathbb{E}_{\bm{g}}\left[\phi_{\bm{g}}\right]$ is itself a convex $G$-Lipschitz function. The claim now follows from Davis & Drusvyatskiy (2019, Theorem 4.4). ∎
@@
In the context of Proposition˜B.5, assume that $\mathcal{D}=\|\cdot\|_{1}$ and that $\mathcal{G}$ is a sum of Dirac measures $\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\bm{g}^{(i)}}$. Then we have further
@@
To obtain the desired bound, and given that we already have the conclusion of Proposition˜B.5, it is enough for us to show that there exists $\mu>0$ such that
$(\forall\bm{m}\in\mathbb{R}^{d})\quad\mu\operatorname{dist}(\bm{m},{\rm{argmin}}~\phi)\leq\phi(\bm{m})-\inf\phi.$
It is a standard result from variational analysis that this property [^6] It is sometimes referred to as (weak) sharp minima (Ferris, 1991; Burke & Ferris, 1993), (superlinear)(linear) conditioning (Lemaire, 1992; Cornejo et al., 1997; Lemaire, 1998), or error bound (Lewis & Pang, 1998). is true if and only if the following inequality is satisfied (see Theorem 5.1 in Cornejo et al. (1997) or Proposition 3.1 in Lemaire (1998))
$(\forall\bm{m}\in\operatorname{dom}\partial\phi)\quad\|\partial\phi(\bm{m})\|_{-}\geq\mu.$
Our goal now is to prove the above inequality, from which the conclusion will follow. Let us consider $\bm{m}\notin{\rm{argmin}}~\phi$, and use standard calculus to write
$\partial\phi(\bm{m})=\frac{1}{n}\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}-{\bm{g}}_{i})=\frac{1}{n}\sum_{i=1}^{n}\big(\operatorname{sgn}(\bm{m}_{j}-({\bm{g}}_{i})_{j})\big)_{j=1,\ldots,d}=\frac{1}{n}\Big(\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}_{j}-({\bm{g}}_{i})_{j})\Big)_{j=1,\ldots,d}.$
Our goal is to show that $\|\partial\phi(\bm{m})\|_{-}\geq\frac{1}{n}$, which is equivalent to show that $\|\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}-{\bm{g}}_{i})\|_{-}\geq 1$. Because $\bm{m}\notin{\rm{argmin}}~\phi$, we know that $0\notin\partial\phi(\bm{m})$ so there must exist a coordinate $j$ such that the $j$-th coordinate of $\partial\phi(\bm{m})$ does not contain zero, in other words such that $0\notin\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}_{j}-({\bm{g}}_{i})_{j})$. Using the fact that $\|\cdot\|_{2}\geq\|\cdot\|_{\infty}$, we can therefore lower bound $\|\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}-{\bm{g}}_{i})\|_{-}\geq|\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}_{j}-({\bm{g}}_{i})_{j})|_{-}$. Let us denote $s_{i}:=\operatorname{sgn}(\bm{m}_{j}-({\bm{g}}_{i})_{j})$, so that we want to prove that $|\sum_{i=1}^{n}s_{i}|_{-}\geq 1$. Let us now consider a few cases.
@@
    where the latter is the distance from the number $s$ to the interval [k,+k][-k,+k]. Now remember that $0\notin\sum_{i=1}^{n}s_{i}$, so this distance cannot be zero, which means that $|s|>k$. Moreover both $s$ and $k$ are rational numbers, so this distance must be greater or equal to $1$. This concludes the proof.
@@ Appendix C Supplementary Information on Experiments
For the language modeling experiments, all details are identical to Kunstner et al. (2023), Section A.1. They also provide implementations for all tasks at https://github.com/fKunstner/noise-sgd-adam-sign, which we use. For the training runs for language modeling, we use batch size $256$ for PTB, $320$ for WikiText-2, and $32$ for SQuAD.
@@ C.1 Additional Plots
Ablations for Fig.˜2
. We provide two ablation studies for the experiment on gradient estimation on heavy-tailed synthetic data. First, Fig.˜7 (left) shows that for Gaussian noise ($\alpha=2$), the performance of momentum can surpass V/CClip depending on selection of the step size $\tau$. However, this does not fix the issue of a quickly degrading performance for momentum, when $\alpha$ decreases.
Second, Fig.˜7 (right) shows the performance under skewed heavy-tailed noise; for this, we set the skewness parameter of the $\alpha$-stable distribution to one.
Figure 7: (Left) Same as Fig.˜2 but with $\tau=0.001$. Here, momentum performs best for Gaussian data, but still becomes instable when the noise is more heavy-tailed. (Right) Same as Fig.˜2 but with skewed noise. Here, the median of the noise is no longer zero, which explains the failure of V/CClip.
Figure 8: Least-squares problem (same as (S1)) for Huber with different values of $\mu$ and $\tau$. Setting $\tau=1.0$ performs relatively well across the selected range of values of $\mu$.
Full gradient estimation with fixed weights.
Here, we show the results for estimating the full gradient of a transformer architecture, with weights fixed at initialization. Fig.˜9 shows the results for PTB where we tried three different batch sizes $\{64,256,1024\}$. Fig.˜10 shows the results for WikiText-2.
We remark that in this experiment only, we have used the first order approximation of (7), namely $\bm{m}_{t+1}=\left(1-\tau\right)\bm{m}_{t}+\tau\bm{g}_{t}$. That is, we replace $\frac{\tau}{1+\tau}$ with its first-order Taylor approximation around zero, which is equal to $\tau$. As the value of $\tau$ is small, this has negligible impact on the result: for example, if $\tau=0.01$, then $\frac{\tau}{1+\tau}\approx 0.0099$.
Discussion. From Fig.˜9, we observe two phenomena: in the long run, momentum attains the lowest error for estimating the full batch gradient. However, the initial decrease of the error is much faster for CClip, followed by VClip. This is important when using these estimates within a training setup such as (15), where we only do one iteration of gradient estimation, followed by an update of the weight (and hence a change in the full-batch gradient). Secondly, we observe that the difference in convergence speeds is most pronounced when the batch size increases. Hence, being robust to outliers seems not to be solvable only by increasing the batch size and thus decreasing the noise of the mini-batch gradient. In fact, the contrary seems to be the case. This is similar to the observations made in Kunstner et al. (2023).